18. The function f(x) = 3 (x -1/2)2 + 4 has

A. Maximum value 4        B. minimum value 4         C. maximum value 4/3       D. minimum value 4/3


1
Expert's answer
2021-03-01T07:13:05-0500

"f(x)=3(x-\dfrac{1}{2})^2+4\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space(1)"

Differentiating above equation with respect to x we get

"f^{'}(x)=6(x-\dfrac{1}{2})\times 1+0\space\space\space\space\space\space\space\space\space\space\space\space\space(2)"

Put "f^{'}(x)=0" in above equation

"\Rightarrow 6(x-\dfrac{1}{2})=0"

"\Rightarrow x=\dfrac{1}{2}"

Put "x=\dfrac{1}{2}" in equation (1), we get

"\Rightarrow f(\dfrac{1}{2})=4"

Differentiating eq.(2) with respect to x,

"f^{''}(x)=6>0\space(local\space minima\space condition)"

"\therefore \space x=\dfrac{1}{2}\space gives\space the\space minimum\space value\space of\space the\space function\space f(x)"

f(x) = 4 is the minimum value of the function.